Grade 11Algebra

Let a,b,c € R. a+b+c=4, then find the maximum value of (ab+bc+ac).

Akarsh
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
square a+b+c to get a^2 + b^2 + c^2 + 2(ab+bc+ca) =  4.
 Note that a^2 + b^2 + c^2 >= ab + bc + ca by AM-GM. 
 
thus ab + bc + ca
 
Hence
 
ab + bc + ca = 16/9 + 16/9 + 16/9
 = 48/9
 = 16/3
 
 
 
Regards
Arun (askIITians forum expert)